If a truck traveling to town b from town a, going at a speed of 45km an hour and another truck traveling to town a from town b, is going at a speed of 54km an hour and 20 minutes later they meet each other. How far is the distance between town a and town b?
Answer by stanbon(48558)

If a truck traveling to town b from town a, going at a speed of 45km an hour and another truck traveling to town a from town b, is going at a speed of 54km an hour and 20 minutes later they meet each other. How far is the distance between town a and town b?
----------------------------
a to b DATA:
rate = 45 km/h ; time = 4/3 hrs; distance = r*t = (4/3)45 = 60 km
-----------------------------
b to a DATA:
rate = 54 km/h ; time = 4/3 hrs; distance = r*t = (4/3)54 = 72 km
-----------------------------
Distance from a to b = 60 + 72 = 132 km
============================================
Cheers,
Stan H.